OpenStudy (anonymous):

If f(x) = −4x + 1, find f(−5).

OpenStudy (anonymous):

f(-5)= -4(-5) +1

OpenStudy (akashdeepdeb):

f(x) = -4x + 1 f(a) = -4a + 1 f(b) = -4b + 1 f(c) = -4c + 1 Find this then! :) f(-5) = -4(-5) + 1 Understood? :)

OpenStudy (anonymous):

apply the remainder theorem let x= -5

OpenStudy (anonymous):

is the answer 21

OpenStudy (akashdeepdeb):

Yes. :)

OpenStudy (anonymous):

thanks

OpenStudy (akashdeepdeb):

:)